OpenStudy (anonymous):

solve 4sin^2 x=4cosx+1

OpenStudy (tkhunny):

No plan? This one is NOT an identity. This one is conditional. Could you solve it if it were this? \(4x^{2} = 4x + 1\) -- No trig, here, just algebra.

OpenStudy (anonymous):

sorry, all solutions are within [0,2pi)

OpenStudy (tkhunny):

You didn't answer my question. The restriction you suggest is common.

OpenStudy (anonymous):

i don't quite understand the question you're asking

OpenStudy (tkhunny):

Okay, let's see if I can explain it better... You are walking down the street, minding your own business. An alien life form appears before you. The life form says, via its universal translator, "I will give you an algebra problem. If you cannot solve it, we will destroy your planet." Solve for x, 4x^2 = 4x + 1 Well, will we survive, or not?

OpenStudy (anonymous):

probably not.

OpenStudy (tkhunny):

Seriously? It's a Quadratic Formula at worst. 4x^2 - 4x - 1 = 0 x = [4 + sqrt(16 - 4(4)(-1))]/(2*4) or [4 - sqrt(16 - 4(4)(-1))]/(2*4) Agreed?

OpenStudy (anonymous):

got you. i was approaching the problem completely different so i was confused why my answer didnt make any sense

OpenStudy (anonymous):

how do i relate this back to my original problem ?

OpenStudy (tkhunny):

No worries. It is often a bit of a shock to get an algebra problem in the middle of a trigonometry problem. This is what we have. 4sin^2 x=4cosx+1 Use the Pythagorean identity to transform to cosine only. 4(1 - cos^2(x))=4cosx+1 So far, so good?

OpenStudy (anonymous):

yep

OpenStudy (tkhunny):

Now, rearrange so that it looks like a normal quadratic equation. 4(1 - cos^2(x)) = 4cosx + 1 4 - 4cos^2(x) = 4cosx + 1 4cos^2(x) + 4cosx - 3 = 0 Now, it's not QUITE a normal Quadratic Equation, but it is so close that if we write cos(x) = instead of x = That is the ONLY difference. Can you solve it for cos(x)? It is a Quadratic Formula, at worst. Perhaps it can be factored.

OpenStudy (anonymous):

for this part 4 - 4cos^2(x) = 4cosx + 1, where did the 4 - come from ?

OpenStudy (anonymous):

sorry just realized you multiplied the 1 inside haha

OpenStudy (tkhunny):

Good. Do you see the final structure?

OpenStudy (anonymous):

-4+-sqrt4^2-4(4)(-3)/2*4

OpenStudy (tkhunny):

Really... Did you try to factor it? 4cos^2(x) + 4cosx - 3 = 0 (2cos(x) + 3)(2cos(x) - 1) = 0

OpenStudy (anonymous):

how do you use the quadratic formula after factoring it ?

OpenStudy (tkhunny):

You don't. The quadratic formula is for stubborn things that cannot be factored. You now use the zero property of multiplication. (2cos(x) + 3)(2cos(x) - 1) = 0 2cos(x) + 3 = 0 cos(x) = -3/2 <== Well, that's just silly. Discard it. Do you see why? 2cos(x) - 1 = 0 cos(x) = 1/2 ==> x = pi/3, 5pi/3

OpenStudy (anonymous):

that was a lot easier than suspected. thank you once again :)

OpenStudy (tkhunny):

Once you see that you need to remember your algebra, you are on your way with this one! If you see those aliens, don't let them intimidate you!

OpenStudy (anonymous):

haha you got it